# A unified relative entropy framework for macroscopic limits of Vlasov–Fokker–Planck equations

### v1

@@ 4.2. High-field limit
As in the diffusive regime, we begin by recording uniform estimates from the entropy–dissipation structure and the kinetic energy bound.
@@
$\varepsilon\partial_{t}f^{\varepsilon}+v\cdot\nabla_{x}f^{\varepsilon}-\nabla(-\Delta)^{-\alpha}\rho^{\varepsilon}\cdot\nabla_{\xi}f^{\varepsilon}+\frac{1}{\varepsilon}\xi^{\perp}\cdot\nabla_{\xi}f^{\varepsilon}=\nabla_{\xi}\cdot(\nabla_{\xi}f^{\varepsilon}+\xi f^{\varepsilon}),\quad t>0,\ (x,\xi)\in\mathbb{R}^{2}\times\mathbb{R}^{2}.$ (4.11)
@@
The preceding entropy and energy bounds ensure that the kinetic system remains uniformly controlled in time, despite the presence of the singular drift induced by the $\xi^{\perp}$ term. To pass from these a priori controls to a quantitative stability estimate, we combine the relative entropy framework with modulated interaction energies, in complete analogy with the diffusive and high–field regimes. The key difference is that the skew-symmetric rotation encoded in $\xi^{\perp}$ introduces no dissipation, but its contribution can still be handled thanks to the uniform $W^{1,\infty}$ bound on the effective velocity $u$ and the smallness of the error term ${\rm e}_{\varepsilon}$. This leads to the following estimate.
@@
$\qquad\leq\mathscr{H}[f^{\varepsilon}|M_{\rho,{\rm u_{\varepsilon}}}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+C\varepsilon.$
@@
$\begin{aligned} &\sup_{0\leq t\leq T}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](t)+\mathscr{P}[\rho^{\varepsilon}|\rho](t)\right)\cr&\quad+\frac{1}{4\varepsilon}\int_{0}^{T}\iint_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\frac{1}{f^{\varepsilon}}|\nabla_{\xi}f^{\varepsilon}+(\xi-u^{\varepsilon})f^{\varepsilon}|^{2}dxd\xi dt+\frac{1}{4\varepsilon}\int_{0}^{T}\int_{\mathbb{R}^{2}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u_{\varepsilon}}|^{2}\,dxdt\cr&\qquad\leq C\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)\right)+C\varepsilon\end{aligned}$ (4.13)
@@
Case of well-prepared initial data. From the uniform bound (4.13), we immediately obtain the strong convergence of both $f^{\varepsilon}$ and $\rho^{\varepsilon}$. To refine this into a quantitative estimate for the flux, we claim that
@@
$I_{2}$ $=-\int_{0}^{T}\iint_{\mathbb{R}^{2}\times\mathbb{R}^{2}}f^{\varepsilon}(\xi\otimes\xi-I_{d}):\nabla\phi^{\perp}\,dxd\xi dt+\int_{0}^{T}\int_{\mathbb{R}^{2}}\rho^{\varepsilon}\nabla^{\perp}\cdot\phi\,dxdt$
@@
$\qquad+\int_{0}^{T}\int_{\mathbb{R}^{2}}(\rho^{\varepsilon}-\rho)\nabla^{\perp}\log\rho\cdot\phi\,dxdt$
@@
This concludes the desired quantitative convergence result.
Case of mildly well-prepared initial data. As in the diffusive and high-field limits, we can extend the convergence analysis to the regime of mildly well-prepared data by working in the bounded Lipschitz topology. In this setting, it follows from Lemmas 3.1 and 3.2, together with (4.13), that
$\sup_{0\leq t\leq T}{\rm d}^{2}_{\rm BL}(\rho^{\varepsilon}(t),\rho(t))+\int_{0}^{T}{\rm d}^{2}_{\rm BL}(f^{\varepsilon},M_{\rho,{\rm u_{\varepsilon}}})\,dt$
@@
$\quad\leq C{\rm d}^{2}_{\rm BL}(\rho^{\varepsilon}_{0},\rho_{0})+C\int_{0}^{T}\iint_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\frac{1}{f^{\varepsilon}}|\nabla_{\xi}f^{\varepsilon}+(\xi-u^{\varepsilon})f^{\varepsilon}|^{2}\,dxd\xi dt+C\int_{0}^{T}\int_{\mathbb{R}^{2}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u_{\varepsilon}}|^{2}\,dxdt$
$\quad\leq C{\rm d}^{2}_{\rm BL}(\rho^{\varepsilon}_{0},\rho_{0})+C\varepsilon\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)\right)+C\varepsilon^{2}.$
In parallel with Section 4.1.1, this also yields
${\rm d}^{2}_{\rm BL}(f^{\varepsilon},M_{\rho,0})\leq C{\rm d}^{2}_{\rm BL}(f^{\varepsilon},M_{\rho,{\rm u_{\varepsilon}}})+C\varepsilon^{2},$
and consequently,
$\int_{0}^{T}{\rm d}^{2}_{\rm BL}(f^{\varepsilon},M_{\rho,0})\,dt\leq C{\rm d}^{2}_{\rm BL}(\rho^{\varepsilon}_{0},\rho_{0})+C\varepsilon\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)\right)+C\varepsilon^{2},$

### v2

@@ 4.2. High-field limit
As in the diffusive regime, we begin by recording uniform estimates from the entropy-dissipation structure and the kinetic energy bound.
@@
$\varepsilon\partial_{t}f^{\varepsilon}+\xi\cdot\nabla_{x}f^{\varepsilon}-\nabla(-\Delta)^{-\alpha}\rho^{\varepsilon}\cdot\nabla_{\xi}f^{\varepsilon}+\frac{1}{\varepsilon}\xi^{\perp}\cdot\nabla_{\xi}f^{\varepsilon}=\nabla_{\xi}\cdot(\nabla_{\xi}f^{\varepsilon}+\xi f^{\varepsilon}),\quad t>0,\ (x,\xi)\in\mathbb{R}^{2}\times\mathbb{R}^{2}.$ (4.13)
@@
The preceding entropy and energy bounds ensure that the kinetic system remains uniformly controlled in time, despite the presence of the singular drift induced by the $\xi^{\perp}$ term. To pass from these a priori controls to a quantitative stability estimate, we combine the relative entropy framework with modulated interaction energies, in complete analogy with the diffusive and high–field regimes. The key difference is that the skew-symmetric rotation encoded in $\xi^{\perp}$ introduces no dissipation, but its contribution can still be handled thanks to the uniform $W^{1,\infty}$ bound on the effective velocity $\rm u_{\varepsilon}$} and the smallness of the error term ${\rm e}_{\varepsilon}$. This leads to the following estimate.
@@
$\quad\leq\mathscr{H}[f^{\varepsilon}|M_{\rho,{\rm u_{\varepsilon}}}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+C\varepsilon.$
@@
$\begin{aligned} &\sup_{0\leq t\leq T}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](t)+\mathscr{P}[\rho^{\varepsilon}|\rho](t)\right)\cr&\quad+\frac{1}{4\varepsilon}\int_{0}^{T}\iint_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\frac{1}{f^{\varepsilon}}|\nabla_{\xi}f^{\varepsilon}+(\xi-u^{\varepsilon})f^{\varepsilon}|^{2}dxd\xi dt+\frac{1}{4\varepsilon}\int_{0}^{T}\int_{\mathbb{R}^{2}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u_{\varepsilon}}|^{2}\,dxdt\cr&\quad\leq C\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)\right)+C\varepsilon\end{aligned}$ (4.15)
@@
From the uniform bound (4.15), we immediately obtain the strong convergence of both $f^{\varepsilon}$ and $\rho^{\varepsilon}$. To refine this into a quantitative estimate for the flux, we claim that
@@
$I_{2}$ $=-\int_{0}^{T}\iint_{\mathbb{R}^{2}\times\mathbb{R}^{2}}f^{\varepsilon}(\xi\otimes\xi-\mathbb{I}_{d}):\nabla\phi^{\perp}\,dxd\xi dt+\int_{0}^{T}\int_{\mathbb{R}^{2}}\rho^{\varepsilon}\nabla^{\perp}\cdot\phi\,dxdt$
@@
$\quad+\int_{0}^{T}\int_{\mathbb{R}^{2}}(\rho^{\varepsilon}-\rho)\nabla^{\perp}\log\rho\cdot\phi\,dxdt$
@@
To derive a quantitative estimate in a negative Sobolev topology, we test the equation for $\rho^{\varepsilon}-\rho$ along the characteristic flow generated by the limiting velocity field. This requires uniform Sobolev control of the associated flow map, together with composition estimates for test functions transported by the flow. We thus record the following auxiliary lemma.
Lemma 4.5.
Let $\gamma\geq 4$, and let
$b\in L^{\infty}(0,T;H^{\gamma}(\mathbb{R}^{2};\mathbb{R}^{2})\cap W^{2,\infty}(\mathbb{R}^{2};\mathbb{R}^{2}))$
be a divergence-free vector field. Let $X(s,t;x)$ be the characteristic flow associated with $b$, namely,
$\frac{d}{ds}X(s,t;x)=b(s,X(s,t;x)),\quad X(t,t;x)=x.$
Then, for every $0\leq s,t\leq T$, the map $X(s,t,\cdot)$ is a $C^{1}$-diffeomorphism of $\mathbb{R}^{2}$ satisfying
$X(s,t,\cdot)-{\rm id}\in H^{\gamma}(\mathbb{R}^{2}),\quad\det\nabla_{x}X(s,t;x)=1\quad\text{for all }x\in\mathbb{R}^{2},$
and
$\|X(s,t,\cdot)-{\rm id}\|_{H^{\gamma}}+\|\nabla_{x}X(s,t,\cdot)\|_{W^{1,\infty}}\leq C\left(\gamma,T,\|b\|_{L^{\infty}(0,T;H^{\gamma}\cap W^{2,\infty})}\right).$
Consequently, for every $\varphi\in H^{\gamma}(\mathbb{R}^{2})$,
$\|\varphi\circ X(s,t,\cdot)\|_{H^{\gamma}}\leq C\left(\gamma,T,\|b\|_{L^{\infty}(0,T;H^{\gamma}\cap W^{2,\infty})}\right)\|\varphi\|_{H^{\gamma}},$
and
$\|\nabla^{2}(\varphi\circ X(s,t,\cdot))\|_{L^{\infty}}\leq C\left(\gamma,T,\|b\|_{L^{\infty}(0,T;H^{\gamma}\cap W^{2,\infty})}\right)\|\varphi\|_{H^{\gamma}}.$
Proof.
Since $b\in L^{\infty}(0,T;W^{1,\infty}(\mathbb{R}^{2};\mathbb{R}^{2}))$, the classical ODE theory yields a unique $C^{1}$-flow $X(s,t;\cdot)$ on $\mathbb{R}^{2}$. Moreover, since $\nabla\cdot b=0$, Liouville’s formula implies
$\det\nabla_{x}X(s,t;x)=1.$
Next, since $b\in L^{\infty}(0,T;H^{\gamma}(\mathbb{R}^{2};\mathbb{R}^{2}))$ with $\gamma>2$, the regularity theory for flows of Sobolev vector fields yields
$X(s,t,\cdot)-{\rm id}\in H^{\gamma}(\mathbb{R}^{2})$
for all $0\leq s,t\leq T$, together with the estimate
$\|X(s,t,\cdot)-{\rm id}\|_{H^{\gamma}}\leq C\left(\gamma,T,\|b\|_{L^{\infty}(0,T;H^{\gamma})}\right).$
Differentiating the flow equation in $x$, we obtain
$\partial_{s}\nabla_{x}X(s,t;x)=\nabla b(s,X(s,t;x))\,\nabla_{x}X(s,t;x),$
and hence
$\|\nabla_{x}X(s,t,\cdot)\|_{L^{\infty}}\leq\exp\left(\int_{s}^{t}\|\nabla b(\tau)\|_{L^{\infty}}\,d\tau\right)\leq C\left(T,\|b\|_{L^{\infty}(0,T;W^{1,\infty})}\right).$
Differentiating once more, for each $1\leq i,j,k\leq 2$ we have
$\partial_{s}\partial_{ij}X^{k}(s,t;x)$ $=\sum_{\ell,m=1}^{2}\partial_{\ell m}b^{k}(s,X(s,t;x))\,\partial_{i}X^{\ell}(s,t;x)\,\partial_{j}X^{m}(s,t;x)$
$\quad+\sum_{\ell=1}^{2}\partial_{\ell}b^{k}(s,X(s,t;x))\,\partial_{ij}X^{\ell}(s,t;x).$
Therefore, by Grönwall’s lemma,
$\|\nabla_{x}^{2}X(s,t,\cdot)\|_{L^{\infty}}\leq C\left(T,\|b\|_{L^{\infty}(0,T;W^{2,\infty})}\right),$
and thus
$\|\nabla_{x}X(s,t,\cdot)\|_{W^{1,\infty}}\leq C\left(\gamma,T,\|b\|_{L^{\infty}(0,T;H^{\gamma}\cap W^{2,\infty})}\right).$
Now, since
$X(s,t,\cdot)-{\rm id}\in H^{\gamma}(\mathbb{R}^{2}),\quad\det\nabla_{x}X(s,t;x)=1,$
the composition estimate of [46, Lemma 2.7] implies
$\|\varphi\circ X(s,t,\cdot)\|_{H^{\gamma}}\leq C\left(\gamma,T,\|b\|_{L^{\infty}(0,T;H^{\gamma}\cap W^{2,\infty})}\right)\|\varphi\|_{H^{\gamma}}.$
Finally, by the chain rule, for each $1\leq i,j\leq 2$,
$\partial_{ij}(\varphi\circ X)=\sum_{k,\ell=1}^{2}(\partial_{k\ell}\varphi)(X)\,\partial_{i}X^{k}\,\partial_{j}X^{\ell}+\sum_{k=1}^{2}(\partial_{k}\varphi)(X)\,\partial_{ij}X^{k},$
and hence
$\|\nabla^{2}(\varphi\circ X)\|_{L^{\infty}}$ $\leq C\left(\|\nabla^{2}\varphi\|_{L^{\infty}}\|\nabla X\|_{L^{\infty}}^{2}+\|\nabla\varphi\|_{L^{\infty}}\|\nabla^{2}X\|_{L^{\infty}}\right)$
$\leq C\left(\gamma,T,\|b\|_{L^{\infty}(0,T;H^{\gamma}\cap W^{2,\infty})}\right)\|\varphi\|_{H^{\gamma}},$
where we used the embedding $H^{\gamma}(\mathbb{R}^{2})\hookrightarrow W^{2,\infty}(\mathbb{R}^{2})$. This completes the proof. ∎
In the present setting, we apply Lemma 4.5 with
$b(t,x):=-\nabla^{\perp}(-\Delta)^{-\alpha}\rho(t,x).$
Assume, in addition, that
$\rho\in L^{\infty}\left(0,T;L^{1}(\mathbb{R}^{2})\cap H^{\gamma+1}(\mathbb{R}^{2})\right).$
We claim that, for every $\alpha\in[\frac{1}{2},1)$,
$\|b(t)\|_{H^{\gamma}}+\|b(t)\|_{W^{2,\infty}}\leq C_{\alpha,\gamma}\left(\|\rho(t)\|_{L^{1}}+\|\rho(t)\|_{H^{\gamma+1}}\right)$
for a.e. $t\in[0,T]$. Indeed, by Plancherel’s theorem,
$\|b(t)\|_{H^{\gamma}}^{2}$ $=\int_{\mathbb{R}^{2}}(1+|\eta|^{2})^{\gamma}|\eta|^{2-4\alpha}|\hat{\rho}(t,\eta)|^{2}\,d\eta$
$=\int_{|\eta|\leq 1}(1+|\eta|^{2})^{\gamma}|\eta|^{2-4\alpha}|\hat{\rho}(t,\eta)|^{2}\,d\eta+\int_{|\eta|>1}(1+|\eta|^{2})^{\gamma}|\eta|^{2-4\alpha}|\hat{\rho}(t,\eta)|^{2}\,d\eta.$
For the low-frequency part, using $\|\hat{\rho}(t)\|_{L^{\infty}}\leq\|\rho(t)\|_{L^{1}}$, we get
$\int_{|\eta|\leq 1}(1+|\eta|^{2})^{\gamma}|\eta|^{2-4\alpha}|\hat{\rho}(t,\eta)|^{2}\,d\eta$ $\leq C\|\rho(t)\|_{L^{1}}^{2}\int_{|\eta|\leq 1}|\eta|^{2-4\alpha}\,d\eta\leq C_{\alpha}\|\rho(t)\|_{L^{1}}^{2},$
where the last integral is finite precisely since $\alpha<1$ in two dimensions.
For the high-frequency part, since $\alpha\geq\frac{1}{2}$, we get $2-4\alpha\leq 0$, and thus $|\eta|^{2-4\alpha}\leq 1$ for $|\eta|>1$. Hence, we have
$\int_{|\eta|>1}(1+|\eta|^{2})^{\gamma}|\eta|^{2-4\alpha}|\hat{\rho}(t,\eta)|^{2}\,d\eta$ $\leq\int_{|\eta|>1}(1+|\eta|^{2})^{\gamma}|\hat{\rho}(t,\eta)|^{2}\,d\eta\leq\|\rho(t)\|_{H^{\gamma}}^{2}\leq\|\rho(t)\|_{H^{\gamma+1}}^{2}.$
Combining the above bounds, we infer that
$\|b(t)\|_{H^{\gamma}}\leq C_{\alpha,\gamma}\left(\|\rho(t)\|_{L^{1}}+\|\rho(t)\|_{H^{\gamma+1}}\right).$
Since $\gamma\geq 4$, the Sobolev embedding $H^{\gamma}(\mathbb{R}^{2})\hookrightarrow W^{2,\infty}(\mathbb{R}^{2})$ yields
$\|b(t)\|_{W^{2,\infty}}\leq C_{\gamma}\|b(t)\|_{H^{\gamma}}\leq C_{\alpha,\gamma}\left(\|\rho(t)\|_{L^{1}}+\|\rho(t)\|_{H^{\gamma+1}}\right).$
Hence
$b\in L^{\infty}(0,T;H^{\gamma}(\mathbb{R}^{2};\mathbb{R}^{2})\cap W^{2,\infty}(\mathbb{R}^{2};\mathbb{R}^{2})),$
and Lemma 4.5 applies.
We now derive a quantitative estimate for $\rho^{\varepsilon}-\rho$ in a negative Sobolev topology. To this end, we assume that $\alpha\in[\frac{1}{2},1)$ and fix an integer $\gamma\geq 4$. Since $L^{1}(\mathbb{R}^{2})\subset H^{-s}(\mathbb{R}^{2})$ for every $s>1$, we have
$\rho^{\varepsilon}(t),\,\rho(t)\in L^{1}(\mathbb{R}^{2})\subset H^{-\gamma}(\mathbb{R}^{2}),\quad t\in[0,T].$
Assume moreover that
$\rho\in L^{\infty}(0,T;H^{\gamma+1}(\mathbb{R}^{2})).$
Let $b(t,x):=-\nabla^{\perp}(-\Delta)^{-\alpha}\rho(t,x)$, and let $X(s,t;x)$ be the characteristic flow associated with $b$, namely,
$\frac{d}{ds}X(s,t;x)=b(s,X(s,t;x)),\quad X(t,t;x)=x.$
Since $\nabla\cdot b=0$, the flow is measure-preserving:
$\det\nabla_{x}X(s,t;x)=1.$
Moreover, by Lemma 4.5,
$\|\varphi(X(s,t;\cdot))\|_{H^{\gamma}}+\|\nabla^{2}(\varphi(X(s,t;\cdot)))\|_{L^{\infty}}\leq C\|\varphi\|_{H^{\gamma}}$
for all $\varphi\in H^{\gamma}(\mathbb{R}^{2})$, where $C=C(\gamma,T,\|\rho\|_{L^{\infty}(0,T;H^{\gamma+1})})$.
On the other hand, arguing as in the diffusive case, one checks that $\rho^{\varepsilon}$ satisfies
$\partial_{t}\rho^{\varepsilon}-\nabla\cdot\left(\rho^{\varepsilon}\nabla^{\perp}(-\Delta)^{-\alpha}\rho^{\varepsilon}\right)=-\nabla\cdot F^{\varepsilon}$
in the sense of distributions, where
$F^{\varepsilon}:=\varepsilon\partial_{t}(m^{\varepsilon})^{\perp}-\nabla^{\perp}\!\cdot\!\int_{\mathbb{R}^{2}}(\xi\otimes\xi-\mathbb{I}_{2})f^{\varepsilon}\,d\xi-(m^{\varepsilon})^{\perp}.$
Here, for a matrix field $A=(A_{ij})_{1\leq i,j\leq 2}$, we define
$(\nabla^{\perp}\cdot A)_{j}:=-\partial_{2}A_{2j}+\partial_{1}A_{1j},\quad j=1,2.$
Testing the equation for $\rho^{\varepsilon}-\rho$ against $\varphi(X(s,t;\cdot))$ and integrating along the flow yields
$\int_{\mathbb{R}^{2}}(\rho^{\varepsilon}-\rho)(t,x)\varphi(x)\,dx$ $=\int_{\mathbb{R}^{2}}(\rho_{0}^{\varepsilon}-\rho_{0})(x)\varphi(X(0,t;x))\,dx$
$\quad-\int_{0}^{t}\int_{\mathbb{R}^{2}}\rho^{\varepsilon}\nabla^{\perp}(-\Delta)^{-\alpha}(\rho^{\varepsilon}-\rho)\cdot\nabla[\varphi(X(s,t;x))]\,dxds$
$\quad-\int_{0}^{t}\int_{\mathbb{R}^{2}}F^{\varepsilon}(s,x)\cdot\nabla[\varphi(X(s,t;x))]\,dxds$
$=:J_{1}+J_{2}+J_{3}.$
For the initial term, Lemma 4.5 gives
$|J_{1}|\leq\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\gamma}}\|\varphi(X(0,t;\cdot))\|_{H^{\gamma}}\leq C\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\gamma}}\|\varphi\|_{H^{\gamma}}.$
For $J_{2}$, we decompose
$J_{2}$ $=-\int_{0}^{t}\int_{\mathbb{R}^{2}}(\rho^{\varepsilon}-\rho)\nabla^{\perp}(-\Delta)^{-\alpha}(\rho^{\varepsilon}-\rho)\cdot\nabla[\varphi(X)]\,dxds$
$\quad-\int_{0}^{t}\int_{\mathbb{R}^{2}}\rho\,\nabla^{\perp}(-\Delta)^{-\alpha}(\rho^{\varepsilon}-\rho)\cdot\nabla[\varphi(X)]\,dxds$
$=:J_{21}+J_{22}.$
Using that $\nabla^{\perp}(-\Delta)^{-\alpha}(\rho^{\varepsilon}-\rho)$ is divergence-free, integrating by parts once, and using
$\|\nabla^{2}(\varphi(X))\|_{L^{\infty}}\leq C\|\varphi\|_{H^{\gamma}},$
by Lemma 4.5, we obtain
$|J_{21}|\leq C\|\varphi\|_{H^{\gamma}}\int_{0}^{t}\|\rho^{\varepsilon}-\rho\|_{\dot{H}^{-\alpha}}^{2}\,ds.$
For $J_{22}$, since $\alpha\in[\frac{1}{2},1)$, by using the same argument as in the proof of Lemma 4.5, we find
$\|(-\Delta)^{-\alpha}\nabla^{\perp}\cdot G\|_{H^{\gamma}}^{2}$ $\leq\int_{\mathbb{R}^{2}}(1+|\eta|^{2})^{\gamma}|\eta|^{2-4\alpha}|\hat{G}(\eta)|^{2}\,d\eta$
$\leq C\|G\|_{L^{1}}^{2}\int_{|\eta|\leq 1}|\eta|^{2-4\alpha}\,d\eta+C\int_{|\eta|>1}(1+|\eta|^{2})^{\gamma}|\hat{G}(\eta)|^{2}\,d\eta$
$\leq C(\|G\|_{L^{1}}^{2}+\|G\|_{H^{\gamma}}^{2}).$
Moreover, since $\gamma>1$ in dimension two, $H^{\gamma}(\mathbb{R}^{2})$ is an algebra, and thus
$\|\nabla\rho\,\varphi(X)\|_{L^{1}}+\|\nabla\rho\,\varphi(X)\|_{H^{\gamma}}\leq C\|\rho\|_{H^{\gamma+1}}\|\varphi(X)\|_{H^{\gamma}}.$
Hence, we have
$|J_{22}|$ $=\left|\int_{0}^{t}\int_{\mathbb{R}^{2}}(\rho^{\varepsilon}-\rho)(-\Delta)^{-\alpha}\nabla^{\perp}\cdot[\nabla\rho\,\varphi(X)]\,dxds\right|$
$\leq C\|\rho\|_{L^{\infty}(0,T;H^{\gamma+1})}\|\varphi(X)\|_{L^{\infty}(0,T;H^{\gamma})}\int_{0}^{t}\|\rho^{\varepsilon}-\rho\|_{H^{-\gamma}}\,ds$
$\leq C\|\varphi\|_{H^{\gamma}}\int_{0}^{t}\|\rho^{\varepsilon}-\rho\|_{H^{-\gamma}}\,ds.$
For $J_{3}$, we decompose according to the three components of $F^{\varepsilon}$:
$J_{3}$ $=\varepsilon\int_{0}^{t}\int_{\mathbb{R}^{2}}\partial_{s}m^{\varepsilon}\cdot\nabla^{\perp}[\varphi(X(s,t;x))]\,dxds$
$\quad-\int_{0}^{t}\int_{\mathbb{R}^{2}}\left(\int_{\mathbb{R}^{2}}(\xi\otimes\xi-\mathbb{I}_{2})f^{\varepsilon}\,d\xi\right):\nabla^{\perp}\otimes\nabla[\varphi(X(s,t;x))]\,dxds$
$\quad-\int_{0}^{t}\int_{\mathbb{R}^{2}}m^{\varepsilon}\cdot\nabla^{\perp}[\varphi(X(s,t;x))]\,dxds$
$=:J_{31}+J_{32}+J_{33}.$
Integrating by parts in time, we obtain
$J_{31}$ $=\varepsilon\int_{\mathbb{R}^{2}}m^{\varepsilon}\cdot\nabla^{\perp}\varphi\,dx-\varepsilon\int_{\mathbb{R}^{2}}m_{0}^{\varepsilon}\cdot\nabla^{\perp}[\varphi(X(0,t;x))]\,dx-\varepsilon\int_{0}^{t}\int_{\mathbb{R}^{2}}m^{\varepsilon}\partial_{s}\nabla^{\perp}[\varphi(X(s,t;x))]\,dxds$
$\leq C\varepsilon\|\varphi\|_{H^{\gamma}}\sup_{0\leq t\leq T}\int_{\mathbb{R}^{2}}\left(\rho^{\varepsilon}|u^{\varepsilon}-{\rm u}_{\varepsilon}|+\rho^{\varepsilon}|{\rm u}_{\varepsilon}|\right)\,dx+C\varepsilon\|\varphi\|_{H^{\gamma}}\int_{\mathbb{R}^{2}}\rho_{0}^{\varepsilon}|u_{0}^{\varepsilon}-{\rm u_{0}}_{\varepsilon}|\,dx+C\varepsilon\int_{\mathbb{R}^{2}}\rho_{0}^{\varepsilon}|{\rm u_{0}}_{\varepsilon}|\,dx$
$\leq C\varepsilon\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon\right)^{\frac{1}{2}}\|\varphi\|_{H^{\gamma}}.$
Here we used
$|\partial_{s}\nabla^{\perp}[\varphi(X(s,t;x))]|$ $=\left|\partial_{s}\left(\nabla X(s,t;x)\cdot(\nabla^{\perp}\varphi)(X(s,t;x))\right)\right|$
$\leq C\|\nabla\nabla^{\perp}(-\Delta)^{-\alpha}\rho\|_{L^{\infty}}\|\nabla\varphi\|_{L^{\infty}}+\|\nabla X(s,t;\cdot)\|_{L^{\infty}}\|\nabla^{2}\varphi\|_{L^{\infty}}\|\nabla^{\perp}(-\Delta)^{-\alpha}\rho\|_{L^{\infty}}$
$\leq C\|\varphi\|_{H^{\gamma}}.$
For $J_{32}$, rewriting it by the integration by parts in $\xi$, we estimate
$J_{32}$ $=\int_{0}^{t}\int_{\mathbb{R}^{2}}\left(\int_{\mathbb{R}^{2}}(\xi f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon})\otimes\xi\,d\xi\right):\nabla^{\perp}\otimes\nabla[\varphi(X(s,t;x))]\,dxds$
@@
$\leq\|\varphi\|_{H^{\gamma}}\left(\int_{0}^{t}\iint_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\frac{1}{f^{\varepsilon}}|\nabla_{\xi}f^{\varepsilon}+\xi f^{\varepsilon}|^{2}\,dxd\xi ds\right)^{\frac{1}{2}}\left(\int_{0}^{t}\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}}|\xi|^{2}f^{\varepsilon}\,dxd\xi ds\right)^{\frac{1}{2}}$
$\leq C\varepsilon^{\frac{1}{2}}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon\right)^{\frac{1}{2}}\|\varphi\|_{H^{\gamma}},$
where we used
$\int_{0}^{t}\iint_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\frac{1}{f^{\varepsilon}}|\nabla_{\xi}f^{\varepsilon}+\xi f^{\varepsilon}|^{2}\,dxd\xi ds$
$\quad=\int_{0}^{t}\iint_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\frac{1}{f^{\varepsilon}}|\nabla_{\xi}f^{\varepsilon}+(\xi-u^{\varepsilon})f^{\varepsilon}|^{2}\,dxd\xi ds+\int_{0}^{t}\int_{\mathbb{R}^{2}}\rho^{\varepsilon}|u^{\varepsilon}|^{2}\,dxds$
$\quad\leq\int_{0}^{t}\iint_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\frac{1}{f^{\varepsilon}}|\nabla_{\xi}f^{\varepsilon}+(\xi-u^{\varepsilon})f^{\varepsilon}|^{2}\,dxd\xi ds+2\int_{0}^{t}\int_{\mathbb{R}^{2}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u_{\varepsilon}}|^{2}\,dxds+2\int_{0}^{t}\int_{\mathbb{R}^{2}}\rho^{\varepsilon}|{\rm u_{\varepsilon}}|^{2}\,dxds$
$\quad\leq C\varepsilon\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)\right)+C\varepsilon^{2},$
due to (4.15).
Finally, using (4.15) again, we deduce
$J_{33}$ $\leq C\|\varphi\|_{H^{\gamma}}\int_{0}^{t}\int_{\mathbb{R}^{2}}\left(\rho^{\varepsilon}|u^{\varepsilon}-{\rm u}_{\varepsilon}|+\rho^{\varepsilon}|{\rm u}_{\varepsilon}|\right)\,dxds$
$\leq C\|\varphi\|_{H^{\gamma}}\left(\int_{0}^{t}\int_{\mathbb{R}^{2}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u}_{\varepsilon}|^{2}\,dxds\right)^{\frac{1}{2}}+C\varepsilon\|\varphi\|_{H^{\gamma}}$
$\leq C\varepsilon^{\frac{1}{2}}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon\right)^{\frac{1}{2}}\|\varphi\|_{H^{\gamma}}.$
Thus, we collect all the estimates for $J_{i}$’s to have
$\|(\rho^{\varepsilon}-\rho)(t)\|_{H^{-\gamma}}\leq\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\gamma}}+C\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon\right)+C\int_{0}^{t}\|(\rho^{\varepsilon}-\rho)(s)\|_{H^{-\gamma}}\,ds.$
An application of Grönwall’s lemma therefore yields
$\|(\rho^{\varepsilon}-\rho)(t)\|_{H^{-\gamma}}\leq C\left(\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\gamma}}+\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon\right),$